# Where to find Grokking in LLM Pretraining? Monitor Memorization-to-Generalization without Test

### v2

@@ Abstract
Grokking, i.e., test performance keeps improving long after training loss converged, has been recently witnessed in neural network training, making the mechanism of generalization and other emerging capabilities such as reasoning mysterious. While existing studies of grokking usually train a small model on one or two highly-specific or toy tasks for thousands of epochs, we conduct the first study of grokking on checkpoints during one-pass pretraining of a 7B large language model (LLM), i.e., OLMoE [15]. We compute the training loss on OLMoE’s pretraining data and evaluate generalization on diverse benchmark tasks, including math reasoning, code generation, and commonsense/domain-specific knowledge retrieval tasks.
Our study, for the first time, verifies that grokking still happens in the pretraining of practical, large-scale foundation models, though different data may enter their grokking stages asynchronously. We further demystify grokking’s “emergence of generalization” by investigating the dynamics of LLMs’ internal states. Specifically, we find that training samples’ pathways (i.e., expert choices across layers) evolve from random, instance-specific to more structured and shareable between samples. Also, the complexity of a single sample’s pathway reduces despite the converged loss. These indicate a memorization-to-generalization “knowledge digestion”, providing a mechanistic explanation of the delayed generalization. In the study, we develop two novel metrics to quantify the distance between pathways and the complexity of a single pathway. We demonstrate their capabilities to predict the generalization improvement on diverse downstream tasks. They are efficient, simple to compute and solely dependent on training data. Hence, they have crucial practical value to model pretraining, enabling us to monitor the generalization performance without finetuning and test. Theoretically, we ground our findings in a spectrum analysis of the tangent kernel for one-layer MoE, showing that more structured pathways reduce model complexity and improve the generalization bound.
@@ 1 Introduction
Foundation models, e.g., large language models (LLMs) with Transformer [23] architectures, have demonstrated that long-term training for a general-purpose task such as autoregressive prediction can gain powerful generalization across diverse downstream tasks and domains [3, 22, 2, 27]. Moreover, some essential capabilities, such as reasoning [7, 1] and prompting [25], also emerge during training. However, a counter-intuitive delayed generalization termed “grokking” [18, 16, 21, 13] has been widely observed in training foundation models: the generalization (e.g., measured by test accuracy) still improves or even sharply rises long after training loss converged or plateaued—a widely accepted indicator of overfitting in conventional machine learning that is detrimental to generalization. While a transition from memorization to generalization is empirically observed during grokking [8], it challenges our fundamental understanding of how and why this transition happens with converged training loss and the underlying mechanism of generalization.
@@
Despite recent attempts to explain the generalization induced by grokking, they have been largely confined to small models trained for hundreds to thousands of epochs on algorithm-synthesized small-scale data for one or two specific tasks [18, 11, 16, 14, 9, 4, 19]. Instead, LLMs’ pre-training often takes only one epoch on a large-scale corpus of real data that might cover different domains and contain noise, imbalances, or redundancy [3, 27]. Hence, whether grokking still emerges in the pretraining of large foundation models remains an open question. Since every sample has been only learned once and the pretraining data are diverse, it is unclear whether and when different samples are memorized, and, whether they are memorized at the same time. Moreover, without repeatedly revisiting the same data, the conversion from memorization to generalization is mysterious: Where does the generalization capability of LLMs come from? When does it happen? Does it rely on memorization in advance (i.e., delayed generalization)?
@@
Unfortunately, these questions cannot be answered by the dynamics of training loss since the generalization performance can still vary with a plateaued training loss, as observed in previous studies of grokking [18, 16]. It is also practically expensive to monitor the generalization performance during LLM pretraining, since a pretrained checkpoint has to be finetuned before being evaluated on downstream tasks due to its lack of instruction-following capability and domain expertise, not mentioning potential confounders introduced by the finetuning recipes. Hence, finding or developing a metric that is efficient to compute and accurate to reflect the dynamics of generalization during pretraining can offer both novel theoretical and practical insights on better understanding and improving the pretraining of foundation models.
@@
Main Contributions   In this paper, we investigate the training dynamics and possible grokking in LLM pretraining to seek for a mechanistic interpretation of its emerging generalization on downstream tasks. Our study is mainly conducted on the open-sourced checkpoints and data for pretraining of a practical-scale LLM, OLMoE [15], which is a mixture-of-experts (MoE) composed of 7-billion parameters. For the first time, we verify that grokking still happens during LLM pretraining. However, unlike the previously observed global and synchronous grokking for most data [18, 14, 16], LLM memorizes different domains/groups of training data at different training steps and takes varying amounts of time to start to generalize to their related test data/tasks. In other words, local grokking still occurs on each group of data but the starting time and lasting steps vary across groups. Due to the local grokking, generalization performance is usually unstable during the earlier pretraining stages but starts to steadily improve once sufficient data have been memorized. In addition, such local difference is related to data difficulty: the later memorized (loss converged) data group often takes longer to generalize.
@@
Since the training loss already converged before grokking starts, it is challenging to explain how the continual pretraining during grokking leads to the observed sharp rise of the generalization performance. In order to demystify the mechanism of memorization-to-generalization, we explore hidden states of LLMs and attempt to track their major changes after loss converges. Specifically, we look into the training dynamics of pathways in OLMoE, i.e., the expert choices across layers for each sample. We introduce two metrics to measure the complexity of pathways and how it changes over time: (1) the edit distance between different samples’ pathways; and (2) the consistency of selected experts’ embedding between consecutive layers for single samples.
@@
Empirical analysis on four domains reveal a prominent trend during grokking that can explain the memorization-to-generalization mechanism: despite more data being introduced to the training, the edit distance declines and the consistency increases. This reflects a decrease on the pathway complexity and a smarter memorization that discovers and leverages more shareable knowledge across samples, leading to more efficient encoding of each individual sample. It explains why generalization improves without reducing the training loss: the model keeps finding more generalizable and shareable representations to memorize training data. Additionally, we provide a theoretical explanation relating the pathway complexity with a generalization bound of a one-layer MoE. Practically, the two metrics show strong correlations with the test performance on standard benchmarks after instruction finetuning. Therefore, they provide a more efficient, test-set/finetuning-free tool to monitor generalization during pretraining, which is essential to model developers and practitioners.
This work takes the first step toward bridging the gaps of understanding grokking and memorization-to-generalization mechanism in LLM pretraining, by systematically analyzing the dynamics of pathway complexity in an MoE LLM. Our key findings and contributions can be summarized as:
@@
  • Grokking still occurs during the one-pass pretraining of practical-scale LLMs but it is local and asynchronous for different data groups/domains, unlike global grokking for all data in previous works.
@@
  • Grokking’s memorization-to-generalization mechanism can be explained by the dynamics of LLMs’ internal states such as MoE pathways, whose similarity between training samples and consistency across layers increase during grokking. They indicate a smarter memorization using more shared knowledge across samples and lower complexity per sample leads to better generalization. This is also supported by a theoretical connection between pathway complexity and a generalization bound.
  • The two novel metrics we developed to measure pathway complexity are computed on training data only without requiring any test/validation set or model finetuning, providing a practically efficient tool to monitor the generalization during LLM pretraining.
@@ 3 Grokking (Delayed Generalization) in LLM Pretraining
A central challenge in large-scale pretraining is identifying when and why models begin to generalize. A common belief holds that once training loss plateaus, generalization has been achieved. Yet, studies on small synthetic tasks reveal a different picture: models often first memorize training data with little generalization, followed by an abrupt shift to high test accuracy—highlighting a significant delay between memorization and generalization [18]. Whether this delayed transition also occurs in LLMs trained on real-world data, however, remains an open question.
In this work, we present the first empirical evidence that delayed generalization also occurs during LLM pretraining—and that training loss is an unreliable proxy for generalization, often diverging from it in both timing and scale. We support this claim at two levels: (1) a domain-level view showing that broad training progress does not necessarily translate into test accuracy gains, and (2) a group-level view revealing that generalization consistently lags behind training, particularly for more difficult examples. These results challenge common assumptions and highlight the need for more faithful metrics to track generalization in LLMs.
We conduct our evaluation using OLMoE’s pretraining data as the training distribution, with corresponding out-of-distribution test sets. Full experimental and dataset details are provided in Appendix B and C.
3.1 Domain-Level Analysis: Local Grokking Emerges Asynchronously across Domains
To analyze how generalization unfolds during pretraining, we first isolate a subset of training samples that the model has not only fit correctly, but has also learned stably over time. We refer to these reliably learned examples as grokking data, borrowing the term “grokking” to emphasize the internal transition from memorization to structured, generalizable representations. Identifying such data allows us to move beyond raw loss convergence and instead focus on the accumulation of enduring internal knowledge that may underpin downstream generalization.
Formally, given a training set $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{N}$ and model checkpoints $\{\theta_{t}\}_{t=1}^{T}$ over $T$ training steps, we track each sample’s loss trajectory $\ell_{i}(t)=\ell(x_{i},y_{i};\theta_{t})$ using cross-entropy loss. We say a sample has converged at time $t_{i}^{*}$—its grokking time—if it satisfies:
$\max_{t^{\prime}\in[t,T]}\ell_{i}(t^{\prime})\leq\varepsilon\quad\text{and}\quad|\ell_{i}(t)-\ell_{i}(T)|\leq\delta,$
where $\varepsilon$ is a low-loss threshold and $\delta$ enforces stability. This filtering ensures that the selected samples reflect robust learning rather than transient memorization.
We use grokking data to track training progress over time and relate it to emerging generalization behavior. Specifically, we analyze the dynamics across four distinct domains: arithmetic reasoning, code generation, commonsense question answering (QA), and domain-specific QA. Figure 1 plots, for each domain, the cumulative number of grokking data, alongside test accuracy on the corresponding downstream tasks.
Training and test dynamics differ across domains. Training convergence and test accuracy evolve in distinct, domain-specific ways. In some cases, training samples steadily converge while test accuracy remains flat or random; in others, accuracy rises early and later declines—well before significant convergence occurs. These divergent patterns reveal that neither memorization nor generalization unfolds in a uniform way across tasks.
Figure 1: Converged groups of pretraining samples (accumulated) over training steps and test accuracy on downstream tasks across four domains. It shows the asynchronous memorization of data from four local domains and how it affects the corresponding generalization: each domain’s test accuracy starts to rise sharply after sufficient training samples converge. But the starting time of grokking and the induced gain vary across domains, resulting in domain-level local grokking rather than global grokking. Math and coding tasks require more knowledge accumulation to generalize than commonsense and domain-specific reasoning.
Within each domain, training progress does not translate directly to generalization. Even as the model continues to learn and memorize training examples, test accuracy gains arrive unevenly—often delayed or decoupled from convergence trends. This highlights a central insight: accumulating converged training samples does not reliably predict when generalization occurs, even within the same domain. Relying on convergence metrics alone can therefore misrepresent the onset and scale of generalization, motivating the need for deeper, more informative indicators.
3.2 Group-Level Analysis: Later Grokking & Longer Delay for More Difficult Data
To better understand this decoupling, we analyze generalization at a finer granularity by tracking when individual training samples are learned. Using the grokking times $t_{i}^{*}$ defined above, we extract the set of samples that are eventually learned stably: $\mathcal{D}_{\text{grok}}=\{(x_{i},y_{i})\mid t_{i}^{*}<\infty\}.$ We then group these training samples based on their grokking onset: $\mathcal{D}_{\text{grok}}^{(t)}=\{(x_{i},y_{i})\mid t_{i}^{*}=t\}.$
For the test set, we similarly classify samples, focusing on those that eventually generalized successfully. To ensure a strict separation from the training set, we apply membership inference attacks [26] to eliminate any test samples that may have been exposed during training. For each remaining sample, we track its prediction trajectory throughout training and assign it to a generalization-time group based on the checkpoint at which its prediction becomes consistently accurate.
For a precise analysis of how improvements in training translate to generalization gains in testing, we pair each training group with the most similar test group. Specifically, we compute embeddings for samples using SentenceTransformer [20] and apply the Hungarian algorithm [10] to perform optimal one-to-one matching between training and test groups based on input similarity.
Figure 2: Training loss and test accuracy over pretraining steps of four data groups in the math domain with local grokking. The pretraining data and test samples in each group are selected and matched based on their semantic similarities. We observe local grokking on every group, i.e., delayed rise of test accuracy after training loss plateaued, though their delayed steps and grokking period lengths differ. From left to right, later converged and memorized (harder) samples exhibit systematically longer delays.
Figure 2 illustrates the key finding: across all convergence-defined training groups and their matched test groups, training loss stabilizes well before test accuracy begins to improve. For training groups that converge early, the delay between convergence and test improvement is short, with accuracy rising soon after the loss flattens. For later-converging groups, the delay is substantially longer, often requiring many additional training steps before generalization appears in the corresponding test groups.
This systematic generalization delay shows that the model requires time not only to memorize the training data but also to reorganize its internal representations into forms that support transfer to unseen examples. This delay manifests not only as a temporal gap, but as a qualitative transition: across groups, test accuracy stays near chance level for an extended period before abruptly rising to perfect generalization during the grokking phase. Importantly, the delay is not random: it scales predictably with sample difficulty, suggesting that the complexity of a pattern determines not just how long it takes to learn, but also how long it takes to generalize. Thus, even at the group level, training loss alone remains an unreliable predictor of generalization readiness.
Open Question. Both domain-level and group-level analyses reveal a core insight: generalization does not arise automatically with training loss convergence. In some cases, models may memorize vast quantities of training data without any meaningful transfer to unseen examples. In other cases, they may require long post-convergence periods before generalization appears. This decoupling of memorization and generalization raises a fundamental question for understanding the emergence of capability in large models:
What internal changes occur within the model that enable this sudden emergence of generalization?
Addressing this question is key to advancing diagnostic tools, improving training efficiency, and uncovering the mechanisms that drive learning in large-scale neural networks. In the next section, we analyze the model’s internal dynamics to tackle this challenge directly.

### v3

# Grokking in LLM Pretraining? Monitor Memorization-to-Generalization without Test

@@ Abstract
This paper presents the first study of grokking in practical LLM pretraining. Specifically, we investigate when an LLM memorizes the training data, when its generalization on downstream tasks starts to improve, and what happens if there is a lag between the two. Unlike existing works studying when a small model generalizes to limited and specified tasks during thousands epochs’ training on algorithmic data, we focus on a practical setting for LLMs, i.e., one-epoch pretraining of next-token prediction on a cross-domain, large-scale corpus, and generalization on diverse benchmark tasks covering math/commonsense reasoning, code generation, and domain-specific retrieval. Our study, for the first time, verifies that grokking still emerges in pretraining mixture-of-experts (MoE) LLMs, though different local data groups may enter their grokking stages asynchronously due to the heterogeneity of their distributions and attributions to others. To find a mechanistic interpretation of this local grokking, we investigate the dynamics of training data’s pathways (i.e., expert choices across layers in MoE). Our primary discovery is that the pathways evolve from random, non-smooth across layers, instance-specific to more structured and transferable across samples, despite the converged pretraining loss. This depicts a transition from memorization to generalization. Two novel metrics are developed to quantify these patterns: one computes the pathway similarity between samples, while the other measures the consistency of aggregated experts between subsequent layers for each sample. These training data based metrics induce zero cost but can faithfully track and monitor the generalization of LLMs on downstream tasks, which, in conventional settings, requires costly instruction tuning and benchmark evaluation. We also ground our findings in a theoretical analysis of one-layer MoE, showing that more structured pathways improve the generalization bound.
@@ 1 Introduction
Large language models (LLMs) with Transformer [25] architectures have demonstrated that pretraining for a general-purpose task such as autoregressive prediction on a large corpus can gain powerful generalization across diverse downstream tasks and domains [3, 24, 2, 29]. Capabilities such as prompting [27], reasoning [8, 1], and in-context learning also emerge during training. However, a counter-intuitive “grokking” [20, 18, 23, 15] has been widely observed in training Transformer models: the generalization starts to improve or even sharply rises long after training loss converged or plateaued—an indicator of overfitting in conventional machine learning. While grokking indicates a transition from memorization to generalization [9], it challenges our fundamental understanding of training dynamics and the underlying mechanism of generalization.
@@
Despite recent attempts to explain the generalization induced by grokking, they have been largely confined to small models trained for hundreds to thousands of epochs on algorithm-synthesized small-scale data with generalization to limited and highly specified tasks [20, 13, 18, 16, 10, 4, 21]. Instead, LLMs’ pre-training often takes only one epoch on a web-scale corpus of heterogeneous, cross-domain data that usually contains noise, imbalance, and redundancy [3, 29]. Hence, LLMs’ learning dynamics can vary drastically across different data, leaving whether grokking still emerges in LLM pretraining an open problem. It is also unclear whether and when different data are memorized with converged loss, and, whether they are memorized at the same time. Moreover, without the repeated replay of the same data and a decrease in loss, the conversion from memorizing training tokens to generalizing on downstream tasks remains mysterious: When does the model start to generalize to different tasks? Is memorization a prerequisite? How does generalization performance change after memorization?
@@
Unfortunately, these questions cannot be answered by the dynamics of pretraining loss since the generalization performance can still vary after the loss plateaued, as observed in previous studies of grokking [20, 18]. For LLMs, it is also practically expensive to monitor its generalization performance during pretraining, since the model has to be finetuned to gain instruction-following capability before being evaluated on downstream tasks. Hence, developing an efficient-to-compute metric to accurately track the generalization can offer critical insights on better understanding and improving the pretraining of LLMs.
@@
Main Contributions   In this paper, we investigate the training dynamics and grokking in LLM pretraining for a mechanistic interpretation of its emergent generalization on downstream tasks. Our study is conducted on the open-sourced pretraining checkpoints of a 7B mixture-of-experts (MoE) LLM, OLMoE [17]. While our study mainly focuses on OLMoE, its pretraining and evaluation setups are common among other LLMs. For the first time, we verify the existence of local grokking during LLM pretraining. However, unlike the previously observed global and synchronous grokking for most data [20, 16, 18], LLM memorizes domains/groups of training data at different training steps and takes varying amounts of steps to generalize to related benchmark tasks. The starting time and lasting steps of such local grokking vary across data groups. As a result, generalization performance is usually unstable during the earlier pretraining stages but starts to steadily improve once sufficient data have been memorized. This local difference also reflects data difficulty: the later memorized data group often takes longer to generalize.
@@
Another primary challenge is to explain how the continual pretraining after memorization completed leads to the delayed sharp rise of the generalization performance. To investigate the mechanism of memorization-to-generalization transition, we explore internal states of LLMs track their major changes after loss converges. Specifically, we study how the pathways (the expert choices across layers for training samples) in OLMoE change. We introduce two metrics to measure their complexity how it changes over time: (1) the edit distance between different samples’ pathways; and (2) the consistency of selected experts’ embedding between consecutive layers for each sample.
@@
Our empirical analysis on four domains reveals a prominent trend of the pathway complexity during grokking: the edit distance declines and the consistency increases, despite the converged training loss and more data being learned. This change in pathway complexity reflects a transition to smarter memorization: it keeps discovering more transferable knowledge across samples to encode data more efficiently. It explains why generalization improves with plateaued training loss: the model keeps finding more generalizable and shareable structures to memorize training data. Additionally, we provide a theoretical explanation relating the pathway complexity to a generalization bound of a one-layer MoE. Practically, the two metrics show strong correlations with the evaluation results on standard benchmarks after instruction finetuning. Therefore, they provide a zero-cost, finetuning and benchmark evaluation-free tool to monitor generalization during pretraining, which is critical to LLM developers and practitioners. These discoveries take the first step toward bridging the gaps of understanding grokking and memorization-to-generalization transition in LLM pretraining. Our key findings and contributions can be summarized as:
@@
  • Grokking still occurs during the one-epoch pretraining of practical-scale LLMs, but it is local and asynchronous for different data groups, unlike global grokking for all data in previous works.
@@
  • Grokking’s memorization-to-generalization transition can be explained by the dynamics of pathways in MoE: pathway similarity between samples and consistency across layers increase. A theoretical connection is also built between the pathway complexity and a generalization bound.
  • The two metrics we developed to measure pathway complexity provide a finetuning and evaluation-free tool with almost zero cost to monitor the generalization during LLM pretraining.
@@ 3 Grokking (Delayed Generalization) in LLM Pretraining
While most existing works study grokking of small models on specified algorithmic tasks, investigating similar phenomena in LLM pretraining poses several new unique challenges, which inherently motivate several designs in our experimental settings.
  • Training-test evaluation gap: most LLMs are trained for the next token(s) prediction while evaluated on diverse benchmark tasks, making the conventional comparison between training and test loss/accuracy invalid. Hence, we need to adapt the definition of grokking to the LLM setting.
  • Heterogeneous, web-sourced, cross-domain training data: While previous grokking studies focus on clean training data with the same distribution/format, most LLMs are pretrained on heterogeneous, cross-domain, web-sourced data that contain redundant, contaminated, or imbalanced samples. Hence, their loss may converge at different rates and pose distinct effects on downstream tasks. To address these challenges, we carefully filter and group the data in our analysis.
  • Generalization to diverse downstream tasks: Previous studies of grokking focus on generalization to highly-specified and limited tasks, while practical LLMs are expected to generalize to arbitrary tasks or domains via prompting. Hence, LLMs’ generalization behaviors may vary drastically across different downstream tasks/benchmarks and difficulty levels during pretraining.
  • One-epoch pretraining: Most LLMs’ pretraining only takes one epoch, so the model has been trained only once on each sample, and its training loss may even converge before visiting later data. However, such “memorization” of unseen data is not caused by overfitting or repeated replay. Instead, it reflects the interpolation or composition of learned data, which might be counted as “generalization” in previous works but differs from the emergent generalization of LLMs on new tasks.
3.1 Study Training Dynamics of LLM Pretraining
Memorization is measured by the pretraining objective. Next token prediction (NTP) is the most widely adopted pretraining objective for LLMs. Let the pretraining corpus be $D_{\text{train}}$, it aims to minimize the negative log-likelihood (NLL) of each token $x_{i,j}$ given previous tokens $x_{i,<j}$ in a text sequence $x_{i}=(x_{i,1},\ldots,x_{i,|j|})$ drawn from $D_{\text{train}}$:
$\ell_{i}(\theta)=-\frac{1}{|x_{i}|}\sum_{j=1}^{|x_{i}|}\log p_{\theta}\!\left(x_{i,j}\,\middle|\,x_{i,<j}\right).$ (1)
Since $\ell_{i}(\theta)$ directly measures how the model memorizes each token in $x_{i}$, we will track it in the pretraining process and identify the sample with consistently small and converged loss across checkpoints as memorized. Specifically, for a sample $x_{i}$ with loss $\ell_{i}(\theta_{t})$ at pretraining step $t$, we identify it as memorized since step $t_{i}^{*}$ if $\ell_{i}(\theta_{t})\leq\varepsilon$ and $|\ell_{i}(\theta_{t})-\ell_{i}(\theta_{T})|\leq\delta$ for all $t\geq t_{i}^{*}$, where $\varepsilon$ is a small threshold, $\delta$ enforces stability over time, and $T$ indexes the final pretraining checkpoint.
Generalization is measured on standard benchmarks after instruction tuning. As foundation models with various emergent capabilities, LLMs are expected to generalize to an open set of unseen tasks through simple prompting/instructions. Hence, their generalization performance is usually measured across standard benchmarks defined on diverse downstream tasks. To equip a pretrained LLM $\theta$ with the instruction-following capability, an instruction tuning algorithm $\mathcal{A}_{\text{IT}}$ further finetunes $\theta$ on a dataset $D_{\text{IT}}$. The generalization performance on benchmark-$i$ is then measured by
$\mathcal{G}_{i}=\text{Eval}_{i}(\theta_{IT},D_{i}),\theta_{IT}=\mathcal{A}_{\text{IT}}(\theta,D_{\text{IT}}),$ (2)
where $D_{i}$ is the dataset in benchmark-$i$ and $\text{Eval}_{i}$ defines the corresponding evaluation metric, which varies across benchmarks for different types of tasks. For example, accuracy is usually used in multi-choice QA tasks while BLEU/ROUGE scores are used in summarization and generation tasks. For code generation tasks, the generalization performance is usually measured by unit tests, where a test case is considered correct if the generated code can pass the unit tests.
Grokking refers to delayed generalization improvement after memorization is completed. Following the same spirit of previous works on grokking, we mainly focus on the phenomenon that $\mathcal{G}_{i}$ starts to surge long after $\ell_{i}(\theta)$ has converged/plateaued. We also use the training loss to track memorization as previous works. However, the test loss adopted by previous grokking analysis is no longer an ideal metric for LLM generalization, since it focuses on token-level matching to ground truths while downstream tasks focus on semantic-level matching, and some of them do not provide ground truths (e.g., code generation).
Experimental Setup We equidistantly sample 10 checkpoints from the OLMoE pretraining trajectory. To evaluate memorization, we probe OLMoE on samples drawn directly from its pretraining corpus $D_{\text{train}}$, ensuring fidelity to the model’s original distribution. We evaluate generalization on widely used benchmarks spanning four domains: math, code, commonsense QA, and domain-specific QA (10k pretraining and 10k test samples per domain; see Appendix A). To exclude contaminated benchmark samples from the generalization analysis, we apply the Min-K%++ membership inference method [28] to identify them. Moreover, we mainly focus on the samples whose model predictions become consistently correct before the pretraining ends, and analyze when this generalization emerges and how long it persists. For downstream tasks, we apply lightweight LoRA-based instruction tuning $\mathcal{A}_{\text{IT}}$ on $D_{\text{IT}}$ to equip checkpoints with basic instruction-following capability while preserving pretrained capabilities (Appendix B).
3.2 Asynchronous Memorization, Delayed Generalization, & Local Grokking
Figure 1: Memorized training samples at different pretraining steps and the generalization performance on standard benchmarks for each domain. It shows asynchronous memorization of different data in each domain and a delayed generalization leap after memorizing a certain amount of data.
To understand when memorization unfolds during pretraining and how it affects generalization, Figure 1 contrasts the number of memorized training samples (using the criterion defined above) at each pretraining checkpoint with benchmark performance for each domain. It shows that (1) training samples are memorized at different pretraining steps; (2) generalization typically follows memorization with a lag: benchmark performance starts to achieve stable gains after sufficient data has been memorized. An interesting discovery is that the lag length varies across domains: generalization leap on math and coding tasks requires memorizing many more samples than commonsense and domain-specific QA tasks.
These observations reveal the complex training dynamics of LLMs caused by training data heterogeneity and their uneven attributions to different benchmark tasks. It implies that the widely studied global grokking—a synchronous memorization of most training data and a delayed contemporary generalization leap on most test samples—cannot be found in LLM pretraining. That being said, the memorization of certain data is plausible to impose a long-term, delayed effect on the generalization to specific downstream tasks. To investigate such local dynamic patterns, we group data in each domain by their memorization steps $t_{i}^{*}$ and particularly focus on the early-memorized groups, as they leave more checkpoints after $t_{i}^{*}$ to study the subsequent dynamics of generalization. Accordingly, we group samples in benchmarks by the training steps since when their predictions become consistently correct. To enable an analysis of how memorization drives downstream generalization performance, we pair each training data group with its most relevant benchmark samples via bipartite graph matching (by Hungarian algorithm [11]), with the training-test data similarity defined in the embedding space of SentenceTransformer [22].
Figure 2: Pretraining objective (blue) and benchmark performance (red) across four paired training-test data groups from the math domain. We observe local grokking on each pair: a delayed generalization leap on benchmark tasks after the pretraining objective plateaued. From left to right, more difficult data memorized (converged) later associate with a longer delay towards grokking.
Figure 2 visualizes the memorization and generalization on four groups: all training-test pairs exhibit a delayed generalization leap on benchmark tasks after the pretraining objective stably converged on the training data. Hence, the training objective cannot monitor or predict the generalization performance. Moreover, earlier-converging groups generalize sooner after loss converged, while later-converged groups associate with substantially longer delays. Notably, this delay scales with data difficulty, which affects not only the number of steps taken to memorize the data but also the duration of memorization to generalization transition.
Remaining Question This local grokking indicates that the model’s internal states, except its pretraining loss, might track generalization or reflect the transition better. This raises a fundamental question for understanding the emergence of capability in LLM:
What internal state changes within an LLM signify the emergence of generalization? Investigating this problem can provide critical insights to uncover the mechanisms of memorization-to-generalization transition and monitor the pretraining progress in practice. To this end, we probe LLMs’ internal state dynamics to address the challenge.